# Structure, short-range order, and phase stability of the AlxCrFeCoNi high-entropy alloy: Insights from a perturbative, DFT-based analysis

### v1

@@ III Discussion
The perturbative analysis also enables extraction of a simple, pairwise model of the internal energy of the alloy, suitable for study via lattice-based atomistic simulations. Here the thermodynamics of each alloy are investigated using the nested sampling algorithm, which efficiently samples different atomic arrangements while estimating their individual contributions to the global partition function. This approach provides greater accuracy and detail in describing thermodynamic phase behaviour compared to mean-field methodologies. We find results in good agreement with the earlier linear-response analysis, along with emergent multi-phase behaviour where different forms of AlNi-FeCoCr phase separation dominate or coexist as temperature changes. The advantage of using nested sampling over Markhov chain Monte Carlo is demonstrated by its ability to sample multiple thermodynamically competitive phases simultaneously.
@@ IV.1 The $\mathbf{S^{(2)}}$ theory for multicomponent alloys
The $S^{(2)}$ theory for multicomponent alloys as used in this work represents a perturbative analysis of the response to the CPA reference medium to an applied inhomogeneous chemical perturbation [36, 37]. The approach uses a Landau-type expansion of the free energy of the system, as evaluated within the CPA, about a homogeneous (disordered) reference state to obtain the two-point correlation function, an ASRO parameter, ab initio. This perturbative analysis can be thought of as analogous to density-functional perturbation theory used for obtaining lattice dynamics, e.g. phonon dispersion relations, within DFT calculations [71]. We stress that effects of the response of the electronic structure and the rearrangement of charge due to the applied chemical perturbation are fully included. The methodology is the natural extension of an earlier modelling approach developed for binary alloys [64, 72, 73], the groundings of which lie in statistical physics and in seminal papers on concentration waves authored by Khachaturyan [42] and Gyorffy and Stocks [43]. Full details of the theory, its implementation, and extensive discussion can be found in a range of earlier works [36, 37, 38, 39, 40, 41].
@@
This formalism is, of course, extensible to multicomponent alloys. As an example, consider a B2 ordering in a Heusler compound with composition $A_{2}BC$. (The B2 structure is visualised in Fig. 11.) An ordering where atoms of species $A$ occupy the 8a sites, while atoms of species $B$ and $C$ remain disordered on the 8b sites, would be described by a wavevector $\mathbf{k}=(0,0,1)$ and equivalent (as for a binary B2 ordering) but with a (normalised) change in concentration $\Delta c_{\alpha}=\frac{1}{\sqrt{6}}(2,-1,-1)$.
Figure 11: Examples of ordered structures which can be described by concentration waves. The L12 structure is an ordered structure on the fcc lattice, while the B2 structure is an ordered structure on the bcc lattice. Both are described by wavevectors $\mathbf{k}=(0,0,2\pi/a)$ and equivalent. Structures are described by their Strukturbericht designation, primitive lattice vectors are denoted by dashed lines, and atomic sites are given their Wyckoff labels.
@@ IV.4 Computational Details
We use the all-electron HUTSEPOT code [59] to generate self-consistent, one electron potentials within the KKR formulation [47, 48] of density functional theory (DFT) [46]. Chemical disorder is described within the coherent potential approximation (CPA) [49, 50, 51], while the the paramagnetic state of the alloys is modelled within the disordered local moment (DLM) picture [56, 57, 58]. We perform spin-polarised, scalar-relativistic calculations within the atomic sphere approximation (ASA) [83]. We use an angular momentum cutoff of $l_{\text{max}}=3$ for basis set expansions, a $20\times 20\times 20$ $\mathbf{k}$-point mesh for integrals over the Brillouin zone, and a 24 point semi-circular Gauss-Legendre grid in the complex plane to integrate over valence energies. We use the local spin-density approximation (LSDA) and the Perdew-Wang exchange-correlation functional [84]. The linear response results are obtained from an computational implementation of the theory described in Section IV.1 and discussed in detail in earlier works [36, 37, 38, 39, 40, 41].
Lattice parameters are set to be consistent with their experimental values [17], and are tabulated in the Supplemental Material [60]. For values of $x$ where the lattice type is not observed experimentally (e.g. bcc, $x=0$; fcc, $x=2$) we perform a volume-conserving transformation from fcc to bcc structures (or vice-versa) to obtain a lattice parameter. For intermediate values of $x$ we interpolate linearly between lattice parameters in accordance with Vegard’s law. For $x=0$, the experimental fcc lattice parameter is 3.57 Å, while for $x=2$, the experimental bcc lattice parameter is 2.88 Å [17]. In a previous study, we have found that the results of the $S^{(2)}$ theory are not particularly sensitive to small deviations in lattice parameters around their experimental values [40].
@@ Acknowledgements.
C.D.W. and J.B.S. conceived the approach. J.B.S. led development of the code for the ASRO linear response calculations. C.D.W. and L.B.P. developed the code implementing the nested sampling algorithm applied to the pairwise atomistic Hamiltonian. Self-consistent DFT calculations were performed by C.D.W. The linear response calculations evaluating ASRO ab initio were performed by J.B.S. and C.D.W., with subsequent analysis and recovery of pairwise atomistic model performed by C.D.W. Atomistic simulations and subsequent analysis were performed by G.A.M. The first draft of the manuscript was written by C.D.W. and subsequently input was received from all authors.
@@
The authors declare no competing interests.

### v2

@@ III Discussion
The perturbative analysis also enables extraction of a simple, pairwise model of the internal energy of the alloy, suitable for study via lattice-based atomistic simulations. Here the thermodynamics of each alloy are investigated using the nested sampling algorithm, which efficiently samples different atomic arrangements while estimating their individual contributions to the global partition function. This approach provides greater accuracy and detail in describing thermodynamic phase behaviour compared to mean-field methodologies. We find results in good agreement with the earlier linear-response analysis, along with emergent multi-phase behaviour where different forms of AlNi-FeCoCr phase separation dominate or coexist as temperature changes. The advantage of using nested sampling over Markov chain Monte Carlo is demonstrated by its ability to sample multiple thermodynamically competitive phases simultaneously.
@@ IV.1 The $\mathbf{S^{(2)}}$ theory for multicomponent alloys
The $S^{(2)}$ theory for multicomponent alloys as used in this work represents a perturbative analysis of the response to the CPA reference medium to an applied inhomogeneous chemical perturbation [36, 37]. The approach uses a Landau-type expansion of the free energy of the system, as evaluated within the CPA, about a homogeneous (disordered) reference state to obtain the two-point correlation function, an ASRO parameter, ab initio. This perturbative analysis can be thought of as analogous to density-functional perturbation theory used for obtaining lattice dynamics, e.g. phonon dispersion relations, within DFT calculations [78]. We stress that effects of the response of the electronic structure and the rearrangement of charge due to the applied chemical perturbation are fully included. The methodology is the natural extension of an earlier modelling approach developed for binary alloys [71, 79, 80], the groundings of which lie in statistical physics and in seminal papers on concentration waves authored by Khachaturyan [43] and Gyorffy and Stocks [44]. Full details of the theory, its implementation, and extensive discussion can be found in a range of earlier works [36, 37, 38, 39, 40, 41, 42].
Figure 11: Illustration of a concentration wave modulating partial atomic site occupancies in a toy, 1D, 3-component alloy. The top panel, (a), shows the homogeneous alloy with no perturbation applied, while the bottom panel, (b), shows the alloy with an example applied chemical perturbation. The alloy’s overall (average) concentrations are $c_{\alpha}=(1/2,1/4,1/4)$, with species coloured blue, green, and red respectively. Real-space lattice sites are at positions $R_{i}=ia$ The normalised chemical polarisation of the example applied concentration wave is $\Delta c_{\alpha}/\|\Delta c_{\alpha}\|=(0,1/\sqrt{2},-1\sqrt{2})$, indicating that partial occupancies of blue atoms remain unchanged, while partial occupancies of green and red atoms are changed by equal and opposite amounts. The associated wave vector is $k=\pi/a$.
@@
to describe a chemical fluctuation, where $\mathbf{R}_{i}$ is the lattice vector with corresponding occupancy $c_{i\alpha}$.
@@
This formalism is naturally applicable to multicomponent alloys. As an example, consider a B2 ordering in a Heusler compound with composition $A_{2}BC$. (The B2 structure is visualised in Fig. 12.) An ordering where atoms of species $A$ occupy the 8a sites, while atoms of species $B$ and $C$ remain disordered on the 8b sites, would be described by a wavevector $\mathbf{k}=(0,0,1)$ and equivalent (as for a binary B2 ordering) but with a (normalised) change in concentration $\Delta c_{\alpha}=\frac{1}{\sqrt{6}}(2,-1,-1)$.
Figure 12: Examples of ordered structures which can be described by concentration waves. The L12 structure is an ordered structure on the fcc lattice, while the B2 structure is an ordered structure on the bcc lattice. Both are described by wavevectors $\mathbf{k}=(0,0,2\pi/a)$ and equivalent for their respective lattice types. Structures are described by their Strukturbericht designation, primitive lattice vectors are denoted by dashed lines, and atomic sites are given their Wyckoff labels.
@@ IV.4 Computational Details
We use the all-electron HUTSEPOT code [60] to generate self-consistent, one electron potentials within the KKR formulation [48, 49] of density functional theory (DFT) [47]. Chemical disorder is described within the coherent potential approximation (CPA) [50, 51, 52], while the the paramagnetic state of the alloys is modelled within the disordered local moment (DLM) picture [57, 58, 59]. We perform spin-polarised, scalar-relativistic calculations within the atomic sphere approximation (ASA) [90]. We use an angular momentum cutoff of $l_{\text{max}}=3$ for basis set expansions, a $20\times 20\times 20$ $\mathbf{k}$-point mesh for integrals over the Brillouin zone, and a 24 point semi-circular Gauss-Legendre grid in the complex plane to integrate over valence energies. We use the local spin-density approximation (LSDA) and the Perdew-Wang exchange-correlation functional [91]. The linear response results are obtained from an computational implementation of the theory described in Section IV.1 and discussed in detail in earlier works [36, 37, 38, 39, 40, 41, 42].
Lattice parameters are set to be consistent with their experimental values [17], and are tabulated in the Supplemental Material [61]. For values of $x$ where the lattice type is not observed experimentally (e.g. bcc, $x=0$; fcc, $x=2$) we perform a volume-conserving transformation from fcc to bcc structures (or vice-versa) to obtain a lattice parameter. For intermediate values of $x$ we interpolate linearly between lattice parameters in accordance with Vegard’s law. For $x=0$, the experimental fcc lattice parameter is 3.57 Å, while for $x=2$, the experimental bcc lattice parameter is 2.88 Å [17]. In a previous study, we have found that the results of the $S^{(2)}$ theory are not particularly sensitive to small deviations in lattice parameters around their experimental values [40, 41].
@@
IV.5 Lattice Fourier Transforms
In sections IV.1 and IV.2, lattice Fourier transforms are invoked to switch between real- and reciprocal-space representations of quantities as needed. We define these transformations explicitly below. In all cases, we consider a system with $N$ Bravais lattice sites. The vector for the real-space position of lattice site $i$ is denoted $\mathbf{R}_{i}$. A scalar quantity $f_{i}$ transforms according to
$f(\mathbf{k})=\frac{1}{\sqrt{N}}\sum_{i}e^{-i\mathbf{k}\cdot\mathbf{R}_{i}}f_{i},$ (18)
while the inverse transformation is defined via
$f_{i}=\frac{1}{\sqrt{N}}\sum_{i}e^{i\mathbf{k}\cdot\mathbf{R}_{i}}f(\mathbf{k}).$ (19)
A pairwise real-space quantity $A_{ij}$, such as the effective pair interaction, $V_{i\alpha;j\alpha^{\prime}}$ of Eq. 14, transforms according to
$A(\mathbf{k})=\frac{1}{\sqrt{N}}\sum_{i}e^{-i\mathbf{k}\cdot\mathbf{R}_{i}}A_{i0},$ (20)
where we have assumed translational invariance. The inverse transformation is then defined via
$A_{ij}=\frac{1}{\sqrt{N}}\sum_{i}e^{i\mathbf{k}\cdot(\mathbf{R}_{i}-\mathbf{R}_{j}})A(\mathbf{k}).$ (21)
Note that the underlying lattice types considered in this work (fcc and bcc) have a monatomic basis, and there is therefore no need to include a sum over basis vectors in any of the above transformations.
@@ Acknowledgements.
C.D.W. and J.B.S. conceived of the approach. J.B.S. led development of the code for the ASRO linear response calculations. C.D.W. and L.B.P. developed the code implementing the nested sampling algorithm applied to the pairwise atomistic Hamiltonian. Self-consistent DFT calculations were performed by C.D.W. The linear response calculations evaluating ASRO ab initio were performed by J.B.S. and C.D.W., with subsequent analysis and recovery of pairwise atomistic model performed by C.D.W. Atomistic simulations and subsequent analysis were performed by G.A.M. The first draft of the manuscript was written by C.D.W. and subsequently input was received from all authors.
@@
J.B.S. acts as an associate editor for npj Computational Materials. The other authors declare no competing interests.